OpenStudy (anonymous):

Solve for x: 1 over 3 (2x − 8) = 4. 2 6 10 10 over 3

OpenStudy (jmark):

Given 1/3(2x − 8) = 4 (2x-8)=3*4 2x-8=12 2x=20 x=10 ICSE class 4 maths question paper - http://icse.edurite.com/icse-question-papers/icse-class-4-maths-question-paper.html